OpenStudy (anonymous):

i need someone to explain the answer for me please which of the following are the two factores of the polynomial t^2-10t+24 the answer is T+6 i got (t-4)(t-6) so why is it t+6

pooja195 (pooja195):

thats right

OpenStudy (anonymous):

but why i did not understand why

pooja195 (pooja195):

 now ask what multiplies to 24 BUT also adds to -10

pooja195 (pooja195):

Making sense so far/

pooja195 (pooja195):

?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i want to understand i got -6 and -4 why the answer was +6???

pooja195 (pooja195):

What was the answer?

pooja195 (pooja195):

You have it right

pooja195 (pooja195):

(t-4)(t-6) When you factor this you get t^2-10t+24

OpenStudy (anonymous):

but i got negative why the answer positive

pooja195 (pooja195):

The answer is right You have it correct BECAUSE when you foil it you get the equation you started with

OpenStudy (anonymous):

ok thank u

pooja195 (pooja195):

(t+−4)(t+−6) =(t)(t)+(t)(−6)+(−4)(t)+(−4)(−6) =t2−6t−4t+24 =t2−10t+24

pooja195 (pooja195):

see t2=t^2